# Stability and convergence of multi-product expansion splitting methods with negative weights for semilinear parabolic equations

### v1

@@ 5.1 Stability of $\mathcal{E}(\tau)$
where the last inequality is obtained by the Hölder inequality.
Without loss of generality, we have
@@
By Grönwall’s lemma, we obtain
$\|u(t)\|_{2}\leq\mathrm{e}^{\kappa t}(\left\|u^{0}\right\|_{2}+C_{0}t),$ (5.5)
where $C_{0}=\left|\Omega\right|^{\frac{1}{2}}|f(0)|$.
@@
$\left\|\partial_{x}^{\mu_{1}}u\partial_{x}^{\mu_{2}}u\cdots\partial_{x}^{\mu_{l}}u\right\|_{2}$ $\leq\left\|\partial_{x}^{\mu_{1}}u\right\|_{4}\left\|\partial_{x}^{\mu_{2}}u\right\|_{4}\left\|\partial_{x}^{\mu_{3}}u\right\|_{\infty}\cdots\left\|\partial_{x}^{\mu_{l}}u\right\|_{\infty},$
$\leq\bar{C}\|u\|^{l}_{H^{m-1}}\leq cM_{m-1}^{l},$
@@
Theorem 5.1 (Uniform norm bounds of exact solution).
Suppose that the initial data $u^{0}\in H^{m}_{\text{per}}(\Omega)$ with $m\geq 1$, $f\in C^{m}(\mathbb{R})$, and (1) is satisfied. Let $u(t)$ be the exact solution to (4.1) with periodic boundary conditions corresponding to the initial data $u^{0}$. Then, for given $T\geq 0$, the exact solution $u(t)$ possesses the uniform Sobolev norm bounds:
where $C\geq 0$ depends on $(\|u^{0}\|_{H^{m-1}(\Omega)},d,\kappa,m,T)$.
In the case of $u^{0}\in L^{p}(\Omega)$ with $p\in[1,\infty]$, $u(t)$ also possesses the uniform $L^{p}$-norm bounds:
where $C\geq 0$ depends on $(|\Omega|,T,|f(0)|)$.

### v2

@@ Appendix A Lie derivative
where the last inequality is obtained by the Hölder inequality. Without loss of generality, we have
@@
By Grönwall’s lemma, we obtain $\|u(t)\|_{2}\leq\mathrm{e}^{\kappa t}(\left\|u^{0}\right\|_{2}+C_{0}t),$ where $C_{0}=\left|\Omega\right|^{\frac{1}{2}}|f(0)|$.
@@
$\left\|\partial_{x}^{\mu_{1}}u\partial_{x}^{\mu_{2}}u\cdots\partial_{x}^{\mu_{l}}u\right\|_{2}$ $\leq\left\|\partial_{x}^{\mu_{1}}u\right\|_{4}\left\|\partial_{x}^{\mu_{2}}u\right\|_{4}\left\|\partial_{x}^{\mu_{3}}u\right\|_{\infty}\cdots\left\|\partial_{x}^{\mu_{l}}u\right\|_{\infty}\leq\bar{C}\|u\|^{l}_{H^{m-1}}\leq\bar{C}M_{m-1}^{l},$
@@
Then the proof of (5.1) can be finished as follows.